OpenStudy (anonymous):

Evaluate 4 + x^3 - x, for x=-3

OpenStudy (anonymous):

wait a second 4 + (-3)^3 - (-3) = 4 - 27 + 3 = -20

OpenStudy (anonymous):

4 + (-3)^3 - (-3) 4 - 27 + 3 -20

OpenStudy (anonymous):

Thank yous